chapter 5 Reviews me 5kg 53 Lorison Jou blog Me: 0.2 2 ی یار 24 the system in equilibrine, ust 06 mas lokg ma: 5kg fs=1 A 6 nI need 3 and 4 answered please
Answer #1

3)

a) If the system is in equilibriu net froce acting on each object must be zero.

Net forcea cting on mB, Fnet = T - mB*g

0 = T - mB*g

T = mB*g

= 5*9.8

= 49 N <<<<<<<<<---------------Answer

as the system is in equilibrium,
a = 0 <<<<<<<<<---------------Answer

Net froce acting on mA, Fnet = T - fs

0 = T - fs

==> fs = T

= 49 N <<<<<<<<<---------------Answer

fs_max = mue_s*NA

= mue_s*mA*g

= 0.6*10*9.8

= 58.8 N <<<<<<<<<---------------Answer


4)
a) if mue_k = 0
Let T is the tention in the string and a is the acceleration of the blocks.

Net force acting on mB, Fnet = mB*g - T

mB*a = mB*g - T

T = mB*g - mB*a ------(1)

Net force acting on mA, Fnet = T

mA*a = T

mA*a = mB*g - mB*a

a*(mA + mB) = mB*g

a = mB*g/(mA + mB)

= 5*9.8/(10 + 5)

= 3.27 m/s^2 <<<<<<<-----------Answer

T = mA*a

= 10*3.27

= 32.7 N <<<<<<<-----------Answer


b) a) if mue_k = 0.5
Let T is the tention in the string and a is the acceleration of the blocks.

Net force acting on mB, Fnet = mB*g - T

mB*a = mB*g - T

T = mB*g - mB*a ------(1)

Net force acting on mA, Fnet = T - fk

mA*a = T - mue_k*mA*g

mA*a = mB*g - mB*a - mue_k*mA*g

a*(mA + mB) = mB*g - mue_k*mA*g

a = (mB*g - mue_k*mA*g)/(mA + mB)

= (5*9.8 - 0.5*10*9.8)/(10 + 5)

= 0 m/s^2 <<<<<<<-----------Answer

T = mB*g - mB*a

= 5*9.8 - 0

= 49 N <<<<<<<-----------Answer